The snow removal business in East Icicle, Minnesota is a competitive industry. All snowplow operators have the cost function C = Q2 + 25; where Q is the number of driveways cleared. Demand for snow removal in the town is given by Qd = 120 - P. The long run equilibrium number of rms in this industry is
(a) 120
(b) 29
(c) 56
(d) 58
(e) 59

(Please show work or explain. I already have the answer. just need to know how to solve it)
Answer #1

In \(\mathrm{LR}, \mathrm{P}=\min \mathrm{AC}=4\left(\mathrm{Q}_{\mathrm{i}}=2\right) .\) Then Qtotal \(=120-4=116\) and \(\mathrm{n}=58\)